# OmniMem: Autoresearch-Guided Discovery of Lifelong Multimodal Agent Memory

### 01 Apr 2026 v1

@@ Abstract
AI agents increasingly operate over extended time horizons, yet their ability to retain, organize, and recall multimodal experiences remains a critical bottleneck. Building effective lifelong memory requires navigating a vast design space spanning architecture, retrieval strategies, prompt engineering, and data pipelines; this space is too large and interconnected for manual exploration or traditional AutoML to explore effectively. We deploy an autonomous research pipeline to discover OmniMem, a unified multimodal memory framework for lifelong AI agents. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, diagnosing failure modes, proposing architectural modifications, and repairing data pipeline bugs, all without human intervention in the inner loop. The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo (0.1170.598) and +214% on Mem-Gallery (0.2540.797) relative to the initial configurations. Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML. We provide a taxonomy of six discovery types and identify four properties that make multimodal memory particularly suited for autoresearch, offering guidance for applying autonomous research pipelines to other AI system domains. Code is available at this https://github.com/aiming-lab/OmniMem.
@@ 1 Introduction
Recent work on autonomous scientific discovery (Lu et al., 2024; Romera-Paredes et al., 2024; Panfilov et al., 2026) has shown that LLM agents can autonomously discover novel algorithms that outperform human-designed baselines, provided the target domain admits well-defined, quantitative evaluation signals. We ask whether this paradigm extends to complex, multi-component AI systems and answer affirmatively. We deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to discover OmniMem, a unified multimodal memory framework for lifelong AI agents. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, iteratively diagnosing failure modes, proposing architectural modifications, repairing data pipeline bugs, and validating improvements, all without human intervention in the inner loop. The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo (0.1170.598) and +214% on Mem-Gallery (0.2540.797) relative to the initial configurations. Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML.
@@
Among the pipeline’s most consequential discoveries are three architectural principles that define OmniMem. First, selective ingestion: lightweight perceptual encoders measure the information novelty of each incoming signal and discard redundant content before storage, significantly reducing storage requirements. Second, unified representation: all memories, regardless of modality, are represented as Multimodal Atomic Units (MAUs) that separate lightweight metadata from heavy raw data, enabling fast search over compact metadata while preserving full-content access on demand. Third, progressive retrieval: a pyramid mechanism expands information in three stages (summaries, details, raw evidence), each gated by a token budget, backed by a hybrid search strategy combining dense vector retrieval with sparse keyword matching via set-union merging, a strategy autonomously discovered by the pipeline. Our key observation is that multimodal memory is particularly well-suited for autonomous research pipelines due to four properties: immediate scalar evaluation metrics enabling tight optimization loops, modular architecture allowing isolated component modification, fast iteration cycles (1–2 hours per experiment) supporting dozens of hypotheses within days, and version-controlled code modifications allowing failed experiments to be cleanly reverted.
@@
In summary, our primary contribution is OmniMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw and that achieves state-of-the-art results on both evaluated benchmarks. Beyond the system itself, we provide a comprehensive taxonomy of autonomous discoveries across experiments, revealing that the highest-impact improvements lie beyond the reach of traditional AutoML, and we characterize the convergence behavior, failure modes, and automatic recovery patterns of the pipeline. Our analysis further identifies four properties that make multimodal memory a particularly suitable domain for autoresearch, providing guidance for future applications to other AI system domains.
@@ 2 Related Work
3 Autoresearch-Guided Discovery of OmniMem
In this section, we describe the autonomous optimization process and the system it produces. We first overview the pipeline (§3.1), then present the discovered OmniMem architecture (§3.2), followed by the benchmark-specific optimization strategy (§3.3).
@@ 3.1 Pipeline Overview
As discussed in Section 1, the design space of multimodal memory systems is too large and interconnected for manual exploration to cover effectively. To address this, we deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to systematically optimize OmniMem. The pipeline receives three inputs: (1) the SimpleMem (Liu et al., 2026a) codebase, a unimodal text-only lifelong memory framework, as a starting point, (2) two benchmark evaluation harnesses with quantitative metrics (F1), and (3) API access to LLM providers. It then enters an iterative loop: at each step, the pipeline analyzes prior results, generates a hypothesis for improvement, implements the change in code, evaluates on a benchmark, and decides whether to proceed (metric improved by ), iterate (ambiguous result; refine the current hypothesis), or pivot (two consecutive degradations; revert and try a new direction). Of total experiments, most resulted in a proceed decision, with the remainder split between iterate and pivot. The full pipeline phases (scoping, literature discovery, multi-agent debate, experiment design, sandboxed execution, analysis, documentation, and finalization) are described in Liu et al. (2026b); here we focus on the discoveries it makes and the mechanisms that enable them.
@@ 3.2.2 Progressive Retrieval with Hybrid Search
Once memories are ingested and stored as MAUs, the next challenge is how to retrieve them efficiently at query time. The second principle is progressive retrieval: rather than loading all retrieved content into the LLM context at once, OmniMem expands information in stages under explicit token budgets.
@@ 3.2.3 Knowledge Graph-Augmented Retrieval
While hybrid search and pyramid retrieval handle queries that can be answered from individual MAUs, many real-world queries require reasoning over multiple connected facts (e.g., “What gift did I give to the person I met at the conference in March?”). The third principle is therefore structured knowledge: OmniMem maintains a knowledge graph that captures entities and relationships across all MAUs.
@@ 4 Experiments
We evaluate OmniMem along two dimensions: (1) the autonomous optimization process, specifically whether the pipeline discovers meaningful improvements across diverse benchmarks, and (2) final system quality, examining whether the discovered architecture achieves state-of-the-art results and whether individual components contribute meaningfully.
@@ 4.3 Main Results
To contextualize these results against existing memory systems, we conduct a controlled comparison of OmniMem against six baselines across five LLM backbones (GPT-4o, GPT-4o-mini, GPT-4.1-nano, GPT-5.1, and GPT-5-nano). Table 1 reports per-category F1 on LoCoMo and five evaluation metrics on Mem-Gallery across all backbones.
On LoCoMo, OmniMem achieves the highest overall F1 across all backbones, ranging from 0.492 (GPT-4.1-nano) to 0.613 (GPT-5.1), substantially outperforming SimpleMem (0.342–0.432), the current state-of-the-art on LoCoMo (Liu et al., 2026a). OmniMem dominates on multi-hop, single-hop, and open-domain categories, with particularly large margins on open-domain questions.
On Mem-Gallery, OmniMem achieves F1 ranging from 0.749 to 0.810, consistently outperforming all memory baselines by a wide margin. SimpleMem is again the strongest baseline (F1 up to 0.538 with GPT-5.1), but still trails OmniMem by over 25 percentage points. These patterns confirm that OmniMem’s gains come from its architectural design (hybrid search, pyramid retrieval, knowledge graph augmentation) rather than from a single dominant component.
@@ 4.4.2 Efficiency
OmniMem achieves 5.81 queries/sec with 8 parallel workers (3.5 faster than the fastest baseline), enabled by read-only FAISS and BM25 indices supporting concurrent lookup (Figure 4, Table 3). All baselines are bottlenecked by sequential LLM generation (85–97% of per-query time), while OmniMem parallelizes the retrieval-generation pipeline via thread-safe read-only indices.
@@
| OmniMem (w=1) | 1.05 | 118 | 846 |
| OmniMem (w=8) | 5.81 | 461 | 821 |
@@ 4.4.3 Case Study: Multi-Hop Retrieval
We illustrate OmniMem’s retrieval pipeline on a real multi-hop query from LoCoMo that requires synthesizing facts across separate conversation sessions. The query asks: “What subject have Caroline and Melanie both painted?” The gold answer is “sunsets,” but answering correctly requires retrieving each person’s painting history from different sessions and identifying the overlap.
Hybrid search. Dense retrieval returns MAUs mentioning Caroline’s paintings (e.g., “Caroline painted a sunset scene”) and Melanie’s art projects (e.g., “Mel and her kids painted a sunset with a palm tree”), but these appear in separate sessions with different surrounding context. BM25 recovers additional MAUs containing the keyword “paint” that rank lower in dense results. Set-union merging preserves the dense ordering and appends BM25-only matches.
@@ 5 Conclusion
We have presented OmniMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw, an autonomous research pipeline. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executed experiments in 72 hours, achieving state-of-the-art on both LoCoMo and Mem-Gallery. The highest-impact discoveries, including bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories), require code comprehension and cross-component reasoning beyond the reach of traditional AutoML. Our taxonomy of six discovery types, together with the observation that multimodal memory is well-suited for autoresearch due to its scalar metrics, modular architecture, and fast iteration cycles, provides a roadmap for applying autonomous research pipelines to other complex AI system domains.
Ethics Statement
OmniMem persistently retains multimodal personal information (text, images, audio, video) over extended time horizons, raising important considerations regarding data privacy and consent. Real-world deployment necessitates explicit user consent, transparent retention policies, and robust mechanisms to review, modify, or delete stored memories. The knowledge graph component, which extracts and links entities across memory units, could enable profiling if deployed without adequate access controls and encryption. Regarding the autonomous research pipeline, all AutoResearchClaw experiments are conducted on established academic benchmarks with no real user data; deploying such pipelines on production systems would require additional oversight. We advocate for memory-augmented agents that prioritize user autonomy, data minimization, and the right to be forgotten.
Reproducibility Statement
We provide complete hyperparameters (Table 4), per-benchmark configurations (Table 5), and system architecture details in Appendix C. Full optimization trajectories with per-iteration metrics and decisions are documented in Appendix E. All prompts are reproduced verbatim in Appendix G, and the AutoResearchClaw pipeline stages are described in Appendix D. Both benchmarks, LoCoMo (Maharana et al., 2024) and Mem-Gallery (Bei et al., 2026), are publicly available; we evaluate under identical splits and protocols as prior work using standard metrics (token-level F1, exact match, BLEU). We open-source the complete framework, benchmark harnesses, and full experiment logs to enable independent replication.
@@ Appendix B Baseline Descriptions
Appendix C OmniMem Technical Details
@@ C.1 System Architecture
The complete OmniMem implementation comprises 13,300 lines of Python across 11 subpackages:
@@ Appendix G Prompt Catalog
This section documents representative prompts used throughout OmniMem, organized by system component. All prompts are reproduced from the source code; minor formatting adjustments are made for readability.
@@ G.2 Modality Processor Prompts
OmniMem employs modality-specific prompts during ingestion to convert heterogeneous inputs into the unified MAU text summary. A key design principle is prompt minimalism: concise instructions yield shorter, more retrieval-friendly summaries at ingestion time, while detailed prompts are reserved for on-demand expansion during answer generation.

### 02 Apr 2026 v2

# Omni-SimpleMem: Autoresearch-Guided Discovery of Lifelong Multimodal Agent Memory

@@ Abstract
AI agents increasingly operate over extended time horizons, yet their ability to retain, organize, and recall multimodal experiences remains a critical bottleneck. Building effective lifelong memory requires navigating a vast design space spanning architecture, retrieval strategies, prompt engineering, and data pipelines; this space is too large and interconnected for manual exploration or traditional AutoML to explore effectively. We deploy an autonomous research pipeline to discover Omni-SimpleMem, a unified multimodal memory framework for lifelong AI agents. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, diagnosing failure modes, proposing architectural modifications, and repairing data pipeline bugs, all without human intervention in the inner loop. The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo and +214% on Mem-Gallery relative to the initial configurations. Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML. We provide a taxonomy of six discovery types and identify four properties that make multimodal memory particularly suited for autoresearch, offering guidance for applying autonomous research pipelines to other AI system domains. Code is available at this https://github.com/aiming-lab/SimpleMem.
@@ 1 Introduction
Recent work on autonomous scientific discovery (Lu et al., 2024; Romera-Paredes et al., 2024; Panfilov et al., 2026) has shown that LLM agents can autonomously discover novel algorithms that outperform human-designed baselines, provided the target domain admits well-defined, quantitative evaluation signals. We ask whether this paradigm extends to complex, multi-component AI systems and answer affirmatively. We deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to discover Omni-SimpleMem, a unified multimodal memory framework for lifelong AI agents. Starting from a naïve baseline (F1 = 0.117 on LoCoMo), the pipeline autonomously executes experiments across two benchmarks, iteratively diagnosing failure modes, proposing architectural modifications, repairing data pipeline bugs, and validating improvements, all without human intervention in the inner loop. The resulting system achieves state-of-the-art on both benchmarks, improving F1 by +411% on LoCoMo (0.1170.598) and +214% on Mem-Gallery (0.2540.797) relative to the initial configurations. Critically, the most impactful discoveries are not hyperparameter adjustments: bug fixes (+175%), architectural changes (+44%), and prompt engineering (+188% on specific categories) each individually exceed the cumulative contribution of all hyperparameter tuning, demonstrating capabilities fundamentally beyond the reach of traditional AutoML.
@@
Among the pipeline’s most consequential discoveries are three architectural principles that define Omni-SimpleMem. First, selective ingestion: lightweight perceptual encoders measure the information novelty of each incoming signal and discard redundant content before storage, significantly reducing storage requirements. Second, unified representation: all memories, regardless of modality, are represented as Multimodal Atomic Units (MAUs) that separate lightweight metadata from heavy raw data, enabling fast search over compact metadata while preserving full-content access on demand. Third, progressive retrieval: a pyramid mechanism expands information in three stages (summaries, details, raw evidence), each gated by a token budget, backed by a hybrid search strategy combining dense vector retrieval with sparse keyword matching via set-union merging, a strategy autonomously discovered by the pipeline. Our key observation is that multimodal memory is particularly well-suited for autonomous research pipelines due to four properties: immediate scalar evaluation metrics enabling tight optimization loops, modular architecture allowing isolated component modification, fast iteration cycles (1–2 hours per experiment) supporting dozens of hypotheses within days, and version-controlled code modifications allowing failed experiments to be cleanly reverted.
@@
In summary, our primary contribution is Omni-SimpleMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw and that achieves state-of-the-art results on both evaluated benchmarks. Beyond the system itself, we provide a comprehensive taxonomy of autonomous discoveries across experiments, revealing that the highest-impact improvements lie beyond the reach of traditional AutoML, and we characterize the convergence behavior, failure modes, and automatic recovery patterns of the pipeline. Our analysis further identifies four properties that make multimodal memory a particularly suitable domain for autoresearch, providing guidance for future applications to other AI system domains.
@@ 2 Related Work
3 Autoresearch-Guided Discovery of Omni-SimpleMem
In this section, we describe the autonomous optimization process and the system it produces. We first overview the pipeline (§3.1), then present the discovered Omni-SimpleMem architecture (§3.2), followed by the benchmark-specific optimization strategy (§3.3).
@@ 3.1 Pipeline Overview
As discussed in Section 1, the design space of multimodal memory systems is too large and interconnected for manual exploration to cover effectively. To address this, we deploy AutoResearchClaw (Liu et al., 2026b), a 23-stage autonomous research pipeline, to systematically optimize Omni-SimpleMem. The pipeline receives three inputs: (1) the SimpleMem (Liu et al., 2026a) codebase, a unimodal text-only lifelong memory framework, as a starting point, (2) two benchmark evaluation harnesses with quantitative metrics (F1), and (3) API access to LLM providers. It then enters an iterative loop: at each step, the pipeline analyzes prior results, generates a hypothesis for improvement, implements the change in code, evaluates on a benchmark, and decides whether to proceed (metric improved by ), iterate (ambiguous result; refine the current hypothesis), or pivot (two consecutive degradations; revert and try a new direction). Of total experiments, most resulted in a proceed decision, with the remainder split between iterate and pivot. The full pipeline phases (scoping, literature discovery, multi-agent debate, experiment design, sandboxed execution, analysis, documentation, and finalization) are described in Liu et al. (2026b); here we focus on the discoveries it makes and the mechanisms that enable them.
@@ 3.2.2 Progressive Retrieval with Hybrid Search
Once memories are ingested and stored as MAUs, the next challenge is how to retrieve them efficiently at query time. The second principle is progressive retrieval: rather than loading all retrieved content into the LLM context at once, Omni-SimpleMem expands information in stages under explicit token budgets.
@@ 3.2.3 Knowledge Graph-Augmented Retrieval
While hybrid search and pyramid retrieval handle queries that can be answered from individual MAUs, many real-world queries require reasoning over multiple connected facts (e.g., “What gift did I give to the person I met at the conference in March?”). The third principle is therefore structured knowledge: Omni-SimpleMem maintains a knowledge graph that captures entities and relationships across all MAUs.
@@ 4 Experiments
We evaluate Omni-SimpleMem along two dimensions: (1) the autonomous optimization process, specifically whether the pipeline discovers meaningful improvements across diverse benchmarks, and (2) final system quality, examining whether the discovered architecture achieves state-of-the-art results and whether individual components contribute meaningfully.
@@ 4.3 Main Results
To contextualize these results against existing memory systems, we conduct a controlled comparison of Omni-SimpleMem against six baselines across five LLM backbones (GPT-4o, GPT-4o-mini, GPT-4.1-nano, GPT-5.1, and GPT-5-nano). Table 1 reports per-category F1 on LoCoMo and five evaluation metrics on Mem-Gallery across all backbones.
On LoCoMo, Omni-SimpleMem achieves the highest overall F1 across all backbones, ranging from 0.492 (GPT-4.1-nano) to 0.613 (GPT-5.1), substantially outperforming SimpleMem (0.342–0.432), the current state-of-the-art on LoCoMo (Liu et al., 2026a). Omni-SimpleMem dominates on multi-hop, single-hop, and open-domain categories, with particularly large margins on open-domain questions.
On Mem-Gallery, Omni-SimpleMem achieves F1 ranging from 0.749 to 0.810, consistently outperforming all memory baselines by a wide margin. SimpleMem is again the strongest baseline (F1 up to 0.538 with GPT-5.1), but still trails Omni-SimpleMem by over 25 percentage points. These patterns confirm that Omni-SimpleMem’s gains come from its architectural design (hybrid search, pyramid retrieval, knowledge graph augmentation) rather than from a single dominant component.
@@ 4.4.2 Efficiency
Omni-SimpleMem achieves 5.81 queries/sec with 8 parallel workers (3.5 faster than the fastest baseline), enabled by read-only FAISS and BM25 indices supporting concurrent lookup (Figure 4, Table 3). All baselines are bottlenecked by sequential LLM generation (85–97% of per-query time), while Omni-SimpleMem parallelizes the retrieval-generation pipeline via thread-safe read-only indices.
@@
| Omni-SimpleMem (w=1) | 1.05 | 118 | 846 |
| Omni-SimpleMem (w=8) | 5.81 | 461 | 821 |
@@ 4.4.3 Case Study: Multi-Hop Retrieval
We illustrate Omni-SimpleMem’s retrieval pipeline on a real multi-hop query from LoCoMo that requires synthesizing facts across separate conversation sessions. The query asks: “What subject have Caroline and Melanie both painted?” The gold answer is “sunsets,” but answering correctly requires retrieving each person’s painting history from different sessions and identifying the overlap.
Hybrid search. Dense retrieval returns MAUs mentioning Caroline’s paintings (e.g., “Caroline painted a sunset”) and Melanie’s art projects (e.g., “Mel and her kids painted a sunset with a tree”), but these appear in separate sessions with different surrounding context. BM25 recovers additional MAUs containing the keyword “paint” that rank lower in dense results. Set-union merging preserves the dense ordering and appends BM25-only matches.
@@ 5 Conclusion
We have presented Omni-SimpleMem, a unified multimodal memory framework whose architecture and configuration are discovered through AutoResearchClaw, an autonomous research pipeline. Starting from a naïve baseline, the pipeline autonomously executed experiments in 72 hours, achieving state-of-the-art on both LoCoMo and Mem-Gallery. The highest-impact discoveries, including bug fixes, architectural changes, and prompt engineering, require code comprehension and cross-component reasoning beyond the reach of traditional AutoML. Our taxonomy of six discovery types, together with the observation that multimodal memory is well-suited for autoresearch due to its scalar metrics, modular architecture, and fast iteration cycles, provides a roadmap for applying autonomous research pipelines to other complex AI system domains.
@@ Appendix B Baseline Descriptions
Appendix C Omni-SimpleMem Technical Details
@@ C.1 System Architecture
The complete Omni-SimpleMem implementation comprises 13,300 lines of Python across 11 subpackages:
@@ Appendix G Prompt Catalog
This section documents representative prompts used throughout Omni-SimpleMem, organized by system component. All prompts are reproduced from the source code; minor formatting adjustments are made for readability.
@@ G.2 Modality Processor Prompts
Omni-SimpleMem employs modality-specific prompts during ingestion to convert heterogeneous inputs into the unified MAU text summary. A key design principle is prompt minimalism: concise instructions yield shorter, more retrieval-friendly summaries at ingestion time, while detailed prompts are reserved for on-demand expansion during answer generation.